if Marwa has a 1,00,000 bond with a 7% interest , how much will she has in 8 years.


1
Expert's answer
2021-07-26T14:08:04-0400

Given:

Present value = 100,000

Interest rate = 7%

Year =8

Future value "=Present \space value \times(1+Interest \space rate)^{years}"

Future value "= 100,000 \times(1+7\%)^8"

Future value"= 100,000 \times 1.71818618"

Future value "= 171,818.618"